A boy shoves his stuffed toy zebra down a frictionless chute. It starts at a height of 1.21 m above the bottom of the chute with an initial speed of 1.39 m/s. The toy animal emerges horizontally from the bottom of the chute and continues sliding along a horizontal surface with a coefficient of kinetic friction of 0.2470. How far from the bottom of the chute does the toy zebra come to rest? Assume ?=9.81 m/s
Answer #1

Initially, it had potential energy and kinetic energy

finally, it will have zero kinetic and potential energy, but work done by friction will be present

mgh + 1/2mv2 = umgd

gh + 1/2v2 = ugd

9.8 * 1.21 + 1/2 * 1.392 = 0.2470 * 9.8 * d

12.82405 = 2.4206 * d

d = 5.297 m